what are the steps to solving this equation? It involves fractions so I hope I write this out for you correctly:
[(-4.57x-3.61)/2.78]-[(3.11x-2.46)/3.44]=1.77x

Answer by stanbon(75887)

what are the steps to solving this equation? It involves fractions so I hope I write this out for you correctly:
[(-4.57x-3.61)/2.78]-[(3.11x-2.46)/3.44]=1.77x
Multiply thru by 2.78*3.44 to get rid of the denominators:
[3.44(-4.57x-3.61)]-[2.78(3.11x-2.46)]=1.77*2.78*3.44x
Multiply thru by 100 to get rid of some of those decimals:
[344(-4.57x-3.61)]-[278(3.11x-2.46)]=177*2.78*3.44x
Distribute:
344*-4.57x - 344*3.61 - 278*3.11x + 278*2.46 = 177*2.78*3.44x
Multiply thru by 100 to get rid of some of those decimals:
344*-457x - 344*361 - 278*311x + 278*246 = 177*278*3.44x
Get all the x-terms together and solve for "x".
===================================================
Cheers,
Stan H.